OpenStudy (anonymous):

Find an equation of the plane that passes through the point and contains the given line. (3, 4, 5) x = 5t, y = 3 + t, z = 4 - t

OpenStudy (amistre64):

the given line gives you one vector on the plane <5,1,-1>

OpenStudy (anonymous):

a(x-3)+b(y-4)+c(z-5) <5,1,-1><a,b,c>=0

OpenStudy (amistre64):

it also gives you a point on the plane (0,3,4)

OpenStudy (amistre64):

use that point and the point given to create another vector with and cross the vectors to get a "normal" vector to the plane

OpenStudy (amistre64):

gonna have to check my work since its been awhile since I tried crossing vectors :) (0,3,4) moves to (3,4,5) ------ <3,1,1> is the vector between the given points <5,1,-1> is the vector from the line equations < i, j ,k> <3, 1, 1> <5,1,-1> (1.-1 - 1.1)i + (1.5 - 3.-1)j + (3.1 - 5.1)k

OpenStudy (amistre64):

<-2,8,-2> if i did it right; lets check by dotting it to the other 2 <-2,8,-2> < 3, 1, 1> ---------- -6 +8 -2 = 0 mighta got lucky <-2,8,-2> < 5, 1,-1> ---------- -10+8+2 = 0 .... looks good

OpenStudy (amistre64):

cant recall if we "simplify" the vector: -2<1,-4,1> maybe? 1,-4,1 1,-4, 1 3, 1, 1 5,1,-1 ------ ------- 3-4+1 = 0 5-4-1 = 0 still works :)

OpenStudy (amistre64):

so lets take our point and apply the normal vector to it to define the plane: 1(x-3) -4(y-4) +1(z-5) = 0 or maybe x -4y +z +8 = 0

OpenStudy (amistre64):

we should get the same results for the second equation regardless of the point we use ... lets try using the point given in the line stuff (0,3,4) 1(x-0)-4(y-3)+1(z-4) = 0 x +4y +z (+12 -4 ) = 0 x +4y +z +8 = 0 ... appears to be good

OpenStudy (anonymous):

Thanks man

OpenStudy (amistre64):

youre welcome :) I see a typo in that last bit .. should be -4y, but its good ;)